OpenStudy (anonymous):

multiply −3 · (−5) · (−2) · (−1) please solve and show work thnk u =)

OpenStudy (amistre64):

15.3 = 45

OpenStudy (amistre64):

15.2 = 30....

OpenStudy (anonymous):

so are u doing the parenth first

OpenStudy (amistre64):

the () are just there to let you know this is multiplication as opposed to subtraction

OpenStudy (anonymous):

o ok so

OpenStudy (anonymous):

ok so y 15 *3? and 15*2

OpenStudy (anonymous):

n y the negavtive became positive?

OpenStudy (amistre64):

you can do this in a ton of ways... 15.2 = 30 10.3=30 5.6 = 30 that 15.3 was a typo on my part :)

OpenStudy (anonymous):

and where the 5*6 come from?

OpenStudy (amistre64):

multiplication is just addition done alot of times 3(5) = 15 means 5+5+5 ; 5 added together 3 times 3(-5) = -15 means (-5) +(-5) +(-5) = -15 -3(-5) = 15 means to do the opposite of addition or: -(-5) -(-5) -(-5) = +(5)+(5)+(5) = 15

OpenStudy (anonymous):

-3*-5=15 -2*-5=10

OpenStudy (amistre64):

there is no 15.3 ; like i said, that was a typo!

OpenStudy (anonymous):

ok so ca u work this out but explai clearly each step

OpenStudy (amistre64):

-2(-5) means -(-2)-(-2)-(-2)-(-2)-(-2) = 10

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

that is why a neg times a neg = a pos

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

can i show you all the elementary steps tp your problem? nah, ill leave that for you to do ;)

OpenStudy (anonymous):

−3 · (−5) · (−2) · (−1) -3*10*-1?

OpenStudy (anonymous):

is these the first steps

OpenStudy (anonymous):

well, if you need more help in integer, try to add me in skype tutorvilleacademy12, i iwll reach more

OpenStudy (anonymous):

annie pls email me @shenita05@aol.com i dont have a skype but i dot have aol instant messenger

OpenStudy (anonymous):

ok i will try to emsil you shenita